# secReferences

### v1

@@ Appendix B P-AUROC on on MVTec 3D-AD
Table 8: Pixel-level anomaly localization results P-AUROC (%) on MVTec 3D-AD with the best in bold.
This section compares our method with other mainstream multimodal AD methods using the P-AUROC metric, including memory bank-based methods BTF [18] and M3DM [28], KD-based methods AST [24] and MMRD [15], and Crossmodal Feature Mapping CFM [12]. As shown in Tab. 8, our method achieves SOTA performance, reaching 99.4% on the P-AUROC metric. Notably, for certain categories like Dowel, our CRD significantly outperforms previous methods. In most other categories, it consistently maintains competitive or near-optimal performance.
@@ Appendix C Additional Results of Ablation Studies
I-AUROC RGB 97.8 94.0 93.9 58.6 98.5 85.3 91.8 62.1 99.2 86.3 86.8
RGB+Crossmodal 98.7 97.6 94.2 65.7 98.4 88.1 91.6 66.5 97.9 90.3 88.9
Depth 99.1 63.8 97.8 99.5 85.3 73.5 99.4 98.8 98.3 55.2 87.1
Depth+Crossmodal 98.3 67.6 98.1 99.1 86.1 76.6 99.2 99.4 98.1 50.5 87.3
RGB+Depth 99.7 89.8 97.8 97.1 99.9 81.3 99.4 96.8 99.9 77.2 93.9
RGB+Depth+Crossmodal 99.6 96.1 98.2 98.5 99.7 90.1 99.2 97.6 100 83.3 96.2
P-AUROC RGB 99.3 99.5 99.5 98.0 99.5 96.1 99.4 99.4 99.4 99.1 98.9
RGB+Crossmodal 99.4 99.6 99.5 98.5 99.5 95.9 99.4 99.4 99.4 99.2 99.0
Depth 99.0 94.0 99.8 99.0 98.7 78.5 99.8 99.9 99.5 93.7 96.2
Depth+Crossmodal 98.9 93.8 99.8 98.9 99.2 84.8 99.8 99.9 99.5 95.9 97.1
RGB+Depth 99.4 99.3 99.7 99.4 99.6 97.2 99.8 99.9 99.6 99.1 99.3
RGB+Depth+Crossmodal 99.3 99.3 99.8 99.3 99.7 97.4 99.8 99.9 99.7 99.5 99.4
PRO RGB 97.2 98.6 98.3 89.4 97.3 87.3 97.8 97.6 95.5 95.3 95.4
RGB+Crossmodal 97.8 98.9 98.2 92.1 97.3 86.6 97.7 97.7 95.5 95.7 95.8
Depth 97.0 81.3 99.1 97.6 95.4 47.5 99.3 99.6 96.2 75.3 88.8
Depth+Crossmodal 96.7 80.0 99.1 97.5 96.9 51.8 99.4 99.6 96.8 82.9 90.1
RGB+Depth 98.2 98.0 99.0 98.2 98.4 90.2 99.3 99.5 98.0 95.4 97.4
RGB+Depth+Crossmodal 97.7 98.1 99.1 98.2 98.7 91.0 99.4 99.5 98.2 97.4 97.7
I-AP RGB 99.4 98.5 98.6 86.9 99.7 96.0 97.8 86.0 99.7 95.3 95.8
RGB+Crossmodal 99.7 99.4 98.7 88.3 99.7 97.0 97.7 88.7 99.2 97.0 96.5
Depth 99.8 88.9 99.5 99.9 95.7 89.8 99.8 99.7 99.3 79.4 95.2
Depth+Crossmodal 99.6 91.2 99.6 99.8 95.7 92.9 99.8 99.9 99.2 79.0 95.7
RGB+Depth 99.9 97.3 99.5 99.2 100 93.9 99.9 98.9 99.9 91.3 98.0
RGB+Depth+Crossmodal 99.9 99.1 99.6 99.6 99.9 97.1 99.8 99.3 100 93.7 98.8
P-AP RGB 49.2 42.9 31.8 47.6 43.6 23.0 37.8 17.9 52.1 31.4 37.7
RGB+Crossmodal 52.2 44.8 30.8 52.3 49.9 23.0 35.2 21.0 50.5 34.0 39.4
Depth 23.9 8.4 45.7 56.4 21.5 8.9 59.7 59.7 49.2 3.3 33.7
Depth+Crossmodal 20.7 2.6 47.3 54.7 26.8 9.9 64.7 58.5 51.5 4.1 34.1
RGB+Depth 42.3 38.2 41.6 71.4 40.7 26.9 58.8 56.7 53.5 22.2 45.2
RGB+Depth+Crossmodal 27.9 36.7 42.8 70.1 50.1 29.4 62.9 57.4 54.8 22.7 45.5
@@
I-AUROC $\mathcal{M}^{RGB}_{norm}+\mathcal{M}^{Dep}_{norm}$ 99.6 96.1 98.2 98.5 99.7 90.1 99.2 97.6 100 83.3 96.2
@@
$(\mathcal{M}^{RGB}/3)\cdot(\mathcal{M}^{Dep}/3)$ [12] 99.6 89.8 98.4 98.9 99.8 82.4 99.0 96.4 100 78.8 94.3
P-AUROC $\mathcal{M}^{RGB}_{norm}+\mathcal{M}^{Dep}_{norm}$ 99.3 99.3 99.8 99.3 99.7 97.4 99.8 99.9 99.7 99.5 99.4
@@
$(\mathcal{M}^{RGB}/3)\cdot(\mathcal{M}^{Dep}/3)$ [12] 99.4 98.9 99.8 99.3 99.7 97.0 99.8 99.9 99.7 99.5 99.3
@@
$(\mathcal{M}^{RGB}/3)\cdot(\mathcal{M}^{Dep}/3)$ [12] 98.0 96.9 99.1 98.2 98.6 88.9 99.4 99.5 98.1 95.7 97.2
@@
$(\mathcal{M}^{RGB}/3)\cdot(\mathcal{M}^{Dep}/3)$ [12] 99.9 97.3 99.7 99.7 100 94.1 99.6 98.9 100 92.5 98.2
@@
$(\mathcal{M}^{RGB}/3)\cdot(\mathcal{M}^{Dep}/3)$ [12] 37.8 28.6 42.5 72.1 45.7 17.6 60.5 54.3 54.2 21.4 43.5
@@ Appendix D Additional Qualitative Results
Figure 6: Qualitative results on MVTec-3D AD. From top to bottom: the input images, the corresponding depth maps, the output anomaly maps of RGB branch, the output anomaly maps of depth branch, the output anomaly maps of CRD, and the ground truth masks.
@@
This section first provides visualizations of the anomaly maps generated by each branch and the fused results, as shown in Fig. 6. The visualizations reveal two key findings: (1) In the depth branch, the significant variation in depth values along edges introduces abundant edge information, making it challenging for student features to fit teacher features at edge regions, leading to false positives. But this issue is less pronounced in the RGB branch. (2) Some anomalies are modality-specific. For instance, discoloration is only detectable in RGB images, while the absence of chocolate on cookies is only evident in depth maps. Consequently, anomaly maps from individual branches may fail to fully capture all anomalies within a sample. However, the fused results effectively achieve comprehensive anomaly localization.
@@
Figure 7: Qualitative results on MVTec-3D AD. From top to bottom: the examples of Bagel, Cable Cland, Cookie, Dowel, and Potato.
We also provide visualizations of anomaly maps before and after incorporating the crossmodal assistants, as in Fig. 7. Without the crossmodal assistants, anomalies that are less prominent in a specific modality goes undetected in the corresponding branch. However, after introducing the crossmodal assistants, this issue is significantly mitigated, resulting in a noticeable reduction in the anomaly miss rate across all modality branches.
Appendix E Results on Eyecandies
E.1 Eyecandies Dataset
The Eyecandies dataset is a synthetic industrial dataset featuring 10 categories of candy samples in an industrial conveyor scenario. Each sample includes RGB images, depth map, and surface normal map. The training set contains 10k samples, with 1k samples per category. The validation and test sets comprise 1k validation samples and 4k test samples, respectively. In our experiments, we select normal maps as the 3D modality and chose images under consistent lighting conditions as the 2D modality for all samples.
E.2 Quantitative result
This section presents the quantitative results of anomaly detection and localization on Eyecandies in Tab. 14. The primary comparison methods include RGB-D [9] and RGB-cD-n [9], proposed alongside the dataset, memory bank-based methods such as BTF [18] and M3DM [28], KD-based methods like AST [24] and MMRD [15], and Crossmodal Feature Mapping CFM [12]. Our method significantly outperforms all non-KD-based multimodal anomaly detection methods. While it slightly lags behind MMRD [15] in anomaly detection metrics I-AUROC, it consistently achieves higher scores than all other methods including P-AUROC and PRO in anomaly localization.
I-AUROC RGB-D [9] 52.9 86.1 73.9 75.2 59.4 49.8 67.9 65.1 83.8 75.0 68.9
RGB-cD-n [9] 59.6 84.3 81.9 84.6 83.3 55.0 75.0 84.6 94.0 84.8 78.7
M3DM [28] 62.4 95.8 95.8 100 88.6 75.8 94.9 83.6 100 100 89.7
AST [24] 57.4 74.7 74.7 88.9 59.6 61.7 81.6 84.1 98.7 98.7 78.0
MMRD [15] 85.4 100 94.6 99.8 90.8 74.7 96.6 98.4 100 100 94.0
CFM [12] 68.0 93.1 95.2 88.0 86.5 78.2 91.7 84.0 99.8 96.2 88.1
CRD (Ours) 86.7 100 98.1 98.1 87.2 72.8 97.1 93.0 100 99.5 93.3
P-AUROC RGB-D [9] 97.3 92.7 95.8 94.5 92.9 80.6 82.7 97.7 93.1 92.8 92.0
RGB-cD-n [9] 98.0 97.9 98.2 97.8 95.1 85.3 97.1 97.8 98.5 96.7 96.2
M3DM [28] 97.4 98.7 96.2 99.8 96.6 94.1 97.3 98.4 99.6 98.5 97.7
AST [24] 76.3 96.0 91.1 96.9 78.8 83.7 91.8 92.4 98.3 96.8 90.2
MMRD [15] - - - - - - - - - - 98.3
CFM [12] 98.3 98.2 96.4 98.9 94.9 94.6 96.9 98.0 99.5 98.7 97.4
CRD (Ours) 99.5 99.0 99.0 99.5 95.7 94.6 99.6 99.0 99.7 99.5 98.5
PRO M3DM [28] 90.6 92.3 80.3 98.3 85.5 68.8 88.0 90.6 96.6 95.5 88.2
AST [24] 51.4 83.5 71.4 90.5 58.7 59.0 73.6 76.9 91.8 87.8 74.4
MMRD [15] 97.5 97.0 94.2 98.5 91.7 68.0 97.0 94.1 99.0 99.2 93.6
CFM [12] 94.2 90.2 83.1 96.5 87.5 76.2 79.1 91.3 93.9 94.9 88.7
CRD (Ours) 98.0 94.1 96.8 98.1 93.6 77.3 98.1 94.4 98.9 98.1 94.7
Table 14: Anomaly detection and localization results I-AUROC (%), P-AUROC (%), and PRO (%) on Eyecandies with the best in bold.
E.3 Qualitative result
@@
Figure 8: Qualitative results on Eyecandies. From top to bottom: the input images, the corresponding depth maps, the output anomaly maps of RGB branch, the output anomaly maps of depth branch, the output anomaly maps of CRD, and the ground truth masks.

### v2

@@ A.1 P-AUC on MVTec 3D-AD
Table 9: Pixel-level anomaly localization results P-AUC (%) on MVTec 3D-AD with the best in bold.
This section compares our method with other mainstream multimodal AD methods on MVTec 3D-AD dataset using the P-AUC metric, including memory bank-based methods BTF [20] and M3DM [31], KD-based methods AST [26] and MMRD [15], and Crossmodal Feature Mapping CFM [12]. As shown in Tab. 9, our method achieves SOTA performance, reaching 99.4% on the P-AUC metric. Notably, for certain categories like Dowel, our CRD significantly outperforms previous methods. In most other categories, it consistently maintains competitive or near-optimal performance.
A.2 P-AUC on Eyecandies
RGB-D [9] 97.3 92.7 95.8 94.5 92.9 80.6 82.7 97.7 93.1 92.8 92.0
RGB-cD-n [9] 98.0 97.9 98.2 97.8 95.1 85.3 97.1 97.8 98.5 96.7 96.2
M3DM [31] 97.4 98.7 96.2 99.8 96.6 94.1 97.3 98.4 99.6 98.5 97.7
AST [26] 76.3 96.0 91.1 96.9 78.8 83.7 91.8 92.4 98.3 96.8 90.2
MMRD [15] - - - - - - - - - - 98.3
CFM [12] 98.3 98.2 96.4 98.9 94.9 94.6 96.9 98.0 99.5 98.7 97.4
CRD (Ours) 99.5 99.0 99.0 99.5 95.7 94.6 99.6 99.0 99.7 99.5 98.5
Table 10: Pixel-level anomaly localization results P-AUC (%) on Eyecandies with the best in bold.
This section presents the quantitative results of anomaly localization on Eyecandies dataset using the P-AUC metric in Tab. 10. The primary comparison methods include RGB-D [9] and RGB-cD-n [9], proposed alongside the dataset, memory bank-based methods such as BTF [20] and M3DM [31], KD-based methods like AST [26] and MMRD [15], and Crossmodal Feature Mapping CFM [12]. Our method significantly outperforms all non-KD-based multimodal anomaly detection methods, achieving 98.5%.
Appendix B Training Curves: I-AUC, P-AUC, and PRO
(a) I-AUC.
(b) P-AUC.
(c) PRO.
Figure 7: AD performance over 500 training epochs.
We provide training curves of I-AUC, P-AUC, and PRO for the carrot and peach categories from the MVTec 3D-AD dataset. As shown in Fig. 7, all metrics stabilize after approximately 50 epochs, indicating that the model converges efficiently without significant performance fluctuations. This case demonstrates the robustness of our method, as it avoids over-generalization even with extended training. Following the setting of Reverse Distillation [13], we adopt 200 epochs for training to ensure sufficient convergence while maintaining computational efficiency.
@@ Appendix C Additional Results of Ablation Studies
I-AUC 2D 97.8 94.0 93.9 58.6 98.5 85.3 91.8 62.1 99.2 86.3 86.8
2D+Crossmodal 98.7 97.6 94.2 65.7 98.4 88.1 91.6 66.5 97.9 90.3 88.9
3D 99.1 63.8 97.8 99.5 85.3 73.5 99.4 98.8 98.3 55.2 87.1
3D+Crossmodal 98.3 67.6 98.1 99.1 86.1 76.6 99.2 99.4 98.1 50.5 87.3
2D+3D 99.7 89.8 97.8 97.1 99.9 81.3 99.4 96.8 99.9 77.2 93.9
2D+3D+Crossmodal 99.6 96.1 98.2 98.5 99.7 90.1 99.2 97.6 100 83.3 96.2
P-AUC 2D 99.3 99.5 99.5 98.0 99.5 96.1 99.4 99.4 99.4 99.1 98.9
2D+Crossmodal 99.4 99.6 99.5 98.5 99.5 95.9 99.4 99.4 99.4 99.2 99.0
3D 99.0 94.0 99.8 99.0 98.7 78.5 99.8 99.9 99.5 93.7 96.2
3D+Crossmodal 98.9 93.8 99.8 98.9 99.2 84.8 99.8 99.9 99.5 95.9 97.1
2D+3D 99.4 99.3 99.7 99.4 99.6 97.2 99.8 99.9 99.6 99.1 99.3
2D+3D+Crossmodal 99.3 99.3 99.8 99.3 99.7 97.4 99.8 99.9 99.7 99.5 99.4
PRO 2D 97.2 98.6 98.3 89.4 97.3 87.3 97.8 97.6 95.5 95.3 95.4
2D+Crossmodal 97.8 98.9 98.2 92.1 97.3 86.6 97.7 97.7 95.5 95.7 95.8
3D 97.0 81.3 99.1 97.6 95.4 47.5 99.3 99.6 96.2 75.3 88.8
3D+Crossmodal 96.7 80.0 99.1 97.5 96.9 51.8 99.4 99.6 96.8 82.9 90.1
2D+3D 98.2 98.0 99.0 98.2 98.4 90.2 99.3 99.5 98.0 95.4 97.4
2D+3D+Crossmodal 97.7 98.1 99.1 98.2 98.7 91.0 99.4 99.5 98.2 97.4 97.7
I-AP 2D 99.4 98.5 98.6 86.9 99.7 96.0 97.8 86.0 99.7 95.3 95.8
2D+Crossmodal 99.7 99.4 98.7 88.3 99.7 97.0 97.7 88.7 99.2 97.0 96.5
3D 99.8 88.9 99.5 99.9 95.7 89.8 99.8 99.7 99.3 79.4 95.2
3D+Crossmodal 99.6 91.2 99.6 99.8 95.7 92.9 99.8 99.9 99.2 79.0 95.7
2D+3D 99.9 97.3 99.5 99.2 100 93.9 99.9 98.9 99.9 91.3 98.0
2D+3D+Crossmodal 99.9 99.1 99.6 99.6 99.9 97.1 99.8 99.3 100 93.7 98.8
P-AP 2D 49.2 42.9 31.8 47.6 43.6 23.0 37.8 17.9 52.1 31.4 37.7
2D+Crossmodal 52.2 44.8 30.8 52.3 49.9 23.0 35.2 21.0 50.5 34.0 39.4
3D 23.9 8.4 45.7 56.4 21.5 8.9 59.7 59.7 49.2 3.3 33.7
3D+Crossmodal 20.7 2.6 47.3 54.7 26.8 9.9 64.7 58.5 51.5 4.1 34.1
2D+3D 42.3 38.2 41.6 71.4 40.7 26.9 58.8 56.7 53.5 22.2 45.2
2D+3D+Crossmodal 27.9 36.7 42.8 70.1 50.1 29.4 62.9 57.4 54.8 22.7 45.5
@@
I-AUC $\mathcal{M}^{RGB}_{norm}+\mathcal{M}^{Dep}_{norm}$ 99.6 96.1 98.2 98.5 99.7 90.1 99.2 97.6 100 83.3 96.2
@@
$\mathcal{M}^{RGB}\cdot\mathcal{M}^{Dep}$ [12] 99.6 89.8 98.4 98.9 99.8 82.4 99.0 96.4 100 78.8 94.3
P-AUC $\mathcal{M}^{RGB}_{norm}+\mathcal{M}^{Dep}_{norm}$ 99.3 99.3 99.8 99.3 99.7 97.4 99.8 99.9 99.7 99.5 99.4
@@
$\mathcal{M}^{RGB}\cdot\mathcal{M}^{Dep}$ [12] 99.4 98.9 99.8 99.3 99.7 97.0 99.8 99.9 99.7 99.5 99.3
@@
$\mathcal{M}^{RGB}\cdot\mathcal{M}^{Dep}$ [12] 98.0 96.9 99.1 98.2 98.6 88.9 99.4 99.5 98.1 95.7 97.2
@@
$\mathcal{M}^{RGB}\cdot\mathcal{M}^{Dep}$ [12] 99.9 97.3 99.7 99.7 100 94.1 99.6 98.9 100 92.5 98.2
@@
$\mathcal{M}^{RGB}\cdot\mathcal{M}^{Dep}$ [12] 37.8 28.6 42.5 72.1 45.7 17.6 60.5 54.3 54.2 21.4 43.5
@@ Appendix D Additional Qualitative Results
Figure 8: Qualitative results on MVTec 3D-AD. From top to bottom: 2D images, 3D depth maps, the output anomaly maps of 2D branch, the output anomaly maps of 3D branch, the output anomaly maps of CRD, and the ground truth masks.
D.1 Qualitative Results on MVTec 3D-AD
@@
This section first provides visualizations of the anomaly maps generated by each branch and the fused results, as shown in Fig. 8. The visualizations reveal two key findings: (1) In the 3D branch, the significant variation in depth values along edges introduces abundant edge information, making it challenging for student features to fit teacher features at edge regions, leading to false positives. But this issue is less pronounced in the 2D branch. (2) Some anomalies are modality-specific. For instance, discoloration is only detectable in RGB images, while the absence of chocolate on cookies is only evident in depth maps. Consequently, anomaly maps from individual branches may fail to fully capture all anomalies within a sample. However, the fused results effectively achieve comprehensive anomaly localization.
@@
Figure 9: Qualitative results on MVTec 3D-AD. From left to right: the examples of Bagel, Cable Cland, Cookie, Dowel, and Potato.
In the main text, we visualize anomaly maps with and without crossmodal assistants, demonstrating their effectiveness in improving anomaly segmentation accuracy by reducing false positives and false negatives. In this section, we provide additional examples to further validate the robustness of our method. As shown in Fig. 9, the inclusion of CF and CA consistently enhances the precision of anomaly localization across various categories, including bagel, cable cland, cookie, dowel, and potato. These results highlight the critical role of crossmodal interactions in refining anomaly detection.
D.2 Qualitative result on Eyecandies
@@
Figure 10: Qualitative results on Eyecandies. From top to bottom: the input 2D images, the corresponding 3D normal maps, the output anomaly maps of 2D branch, the output anomaly maps of 3D branch, the output anomaly maps of CRD, and the ground truth masks.